OpenStudy (anonymous):

Simplify using order of operations: 3^3 - (8 x 3). A. -15 B. -3 C. -2 D. 3

jigglypuff314 (jigglypuff314):

order of operations: PEMDAS Parenthesis Exponents Multiplication/Division (left to right) Addition/Subtraction (left to right) so for 33 - (8 x 3) multiply first then subtract.

OpenStudy (anonymous):

i don't understand this

jigglypuff314 (jigglypuff314):

33 - (8 x 3) you would multiply first so 8 x 3 =

OpenStudy (anonymous):

24

jigglypuff314 (jigglypuff314):

then subtract so 33 - 24 =

OpenStudy (anonymous):

9

jigglypuff314 (jigglypuff314):

yes... but now I'm wondering why it does not match up with any of the options :/

OpenStudy (anonymous):

yeah i was wondering the same thing

jigglypuff314 (jigglypuff314):

are you sure you aren't missing anything in the question? or copied the correct options?

OpenStudy (anonymous):

oh yes sorry about that its 3^3

jigglypuff314 (jigglypuff314):

okay, so what is 3^3 = ?

OpenStudy (anonymous):

9

jigglypuff314 (jigglypuff314):

3^3 = 3 x 3 x 3 = and might I ask for the full question?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

and i don't what you meant by full question

OpenStudy (anonymous):

http://prntscr.com/24x3kk

jigglypuff314 (jigglypuff314):

oh okay, so since we already know that 8 x 3 = 24 then what is 27 - 24 = ?

OpenStudy (anonymous):

3

jigglypuff314 (jigglypuff314):

yep :)